OpenStudy (anonymous):

someone help! find the slope of the line that passes through the points (5,8) and (9,4) A. 3 B. 1 C. -1 D. 4/3

OpenStudy (tamtoan):

do you know the slope formula ?

OpenStudy (anonymous):

I THINK Y=MX+B

OpenStudy (anonymous):

ACTUALLY IDK

OpenStudy (tamtoan):

that is the equation formula with slope m ...you need the slope formula m = (y2 - y1)/(x2 - x1) ...with 2 points given ...(x1,y1) and x2,y2) can you try ?

OpenStudy (anonymous):

here is the slope form x1 = 1, y1 = - 4, x2 = -4, and y2 = 2

OpenStudy (tamtoan):

no, your 2 points are : (5,8) and (9,4)

OpenStudy (anonymous):

ya so dont i take 5-8 amd 8-4

OpenStudy (anonymous):

i mean 4-8/9-5

OpenStudy (anonymous):

so -1 is our answer right?

OpenStudy (tamtoan):

put in parentheses :) or else your computation will not work, can you try again and see what you get ?

OpenStudy (tamtoan):

correct ...but the way you typed there ...it's \[4 - \frac{ 8 }{ 9 } - 5\] and this won't come out -1 :)

OpenStudy (anonymous):

yay!

OpenStudy (anonymous):

(8-4)/(5-9)=-1